Find the general/particular solution of the following Differential Equations.


(Integrable Combinations)


1.) (xdy-ydx)/(x^2) = x^3 dx


1
Expert's answer
2021-09-13T14:36:22-0400

"\frac{{xdy - ydx}}{{{x^2}}} = {x^3}dx \Rightarrow xdy - ydx = {x^5}dx \Rightarrow x\frac{{dy}}{{dx}} - y = {x^5} \Rightarrow \frac{{dy}}{{dx}} - \frac{y}{x} = {x^4} \Rightarrow y' - \frac{y}{x} = {x^4}"

Let

"y = u(x)v(x) \Rightarrow y' = u'v + uv'"

Then

"u'v + uv' - \frac{{uv}}{x} = {x^4} \Rightarrow u'v + u\left( {v' - \frac{v}{x}} \right) = {x^4}"

Let

"v' - \frac{v}{x} = 0 \Rightarrow \frac{{dv}}{{dx}} = \frac{v}{x} \Rightarrow \frac{{dv}}{v} = \frac{{dx}}{x} \Rightarrow \ln v = \ln x \Rightarrow v = x"

Then

"u'x = {x^4} \Rightarrow u' = {x^3} \Rightarrow u = \frac{{{x^4}}}{4} + C"

Then

"y = uv = \left( {\frac{{{x^4}}}{4} + C} \right)x = \frac{{{x^5}}}{4} + Cx"

Answer: "y = \frac{{{x^5}}}{4} + Cx"